OpenStudy (anonymous):

sin (x/2) = -cos (x)

OpenStudy (solomonzelman):

Solving for x?

OpenStudy (anonymous):

how

OpenStudy (anonymous):

sorry but i don't get it

OpenStudy (solomonzelman):

\[Identity:~~~~~~\sin(\frac{x}{2})~=~±\sqrt{\frac{1-\cos(x)}{2}}\]

OpenStudy (anonymous):

ok , then how would you solve for x ?

OpenStudy (anonymous):

I'm stupid, Sorry !!!

OpenStudy (solomonzelman):

lets rewrite it, \[-\cos(x)=\sqrt{\frac{1-Cos(x)}{2}}~~~~~~~~~~~~square~~both~~sides,\]\[Cos^2x=\frac{1-Cos(x)}{2}~~~~~~~~~~~~~~~~~~~`~both~~sides~~times~~2,\]\[2Cos^2x=1-Cos(x)\]\[Let~~~~Cos(x)=a,~~~~Do~~i t.\]

OpenStudy (anonymous):

x is an angle

OpenStudy (solomonzelman):

So it will now be\[2a^2=1-a,~~~~~~add~~a,~~and~~then~~subtract~~1~~~~~~from~~both~~sides.\]

OpenStudy (solomonzelman):

Lets first solve for a, or for cosx, and whet we get for a, substitute into cosx=a

OpenStudy (anonymous):

so what will be the angle ?

OpenStudy (solomonzelman):

Solve for a. I'll do it in my had, I'll use quadratic.

OpenStudy (solomonzelman):